OpenStudy (anonymous):

Together Maria and Johannes can pick a bushel of apples in 16 min. Maria takes 4 min. more to pick a bushel than johannes does. Find the time it takes each person to pick a bushel of apples alone?

OpenStudy (anonymous):

Find the time it takes each person to pick a bushel of apples alone. : Let m = time for Maria to pick a bushel alone then (m-4) = time for Joan to do it : Let the completed job = 1 (a bushel of apples : Each girl will pick a fraction of a bushel and the two fractions must equal 1 \[\frac{16}{m}+\frac{16}{m-4}=1\] Multiply by m(m-4), results: 16(m-4) + 16m = m(m-4) : 16m - 64 + 16m = m^2 - 4m combine like terms on the right 0 = m^2 - 4m - 32m + 64 A quadratic equation m^2 - 36m + 64 = 0 This will not factor, use the quadratic formula: a=1; b=-36; c=64 \[ m=\frac{-(-36)\pm \sqrt{-36^{2}-(4\times 1 \times 4)}}{2\times1}\] here your answer, good luck your home work

OpenStudy (anonymous):

If I deserve medal then give me one

OpenStudy (anonymous):

i thik u desereve 2!